# Debiasing Watermarks for Large Language Models via Maximal Coupling

### v1

@@ 3.2 Detection
Detection as Hypothesis Testing.
Let $m$ be the number of green tokens in $\tilde{w}_{1:n}$. For simplicity, we assume that the first $m$ tokens in $\tilde{w}_{1:n}$, namely $\tilde{w}_{1},...,\tilde{w}_{m}$, represent the green tokens. If no corruption was made to the generated text, we can reformulate the detection objective into the following hypothesis testing problem:
@@ 2. $H_{1}$:
Sparse Mixture Detection.
In practice, the user may modify the generated text. For example, when a token is substituted, the corresponding uniform random variable $\zeta$ would be independent of the variable whether the new token is from the green list. In this case, the statistical signal of the watermark becomes sparse, leading us to consider an alternative hypothesis where a fraction $\varepsilon_{m}$ of the $\zeta_{t}$’s still present the watermark signal. Thus, we propose the following:
@@ 1. $H_{1}^{(\mathrm{mix})}$:
Sum Test.
Since the probability sum of the green tokens, $P_{t,{\mathcal{G}}_{t}}$’s, are unknown to the detector, the likelihood ratio test cannot be used in practice. For this reason, we begin by considering a simple test based on the sum $s=\sum_{i=1}^{m}\zeta_{m}$. Under the null hypothesis, $\mathbb{E}_{0}s=m/2$, while under the alternative hypothesis, $\mathbb{E}_{1}s$ is smaller than $m/2$ by a factor depending on $m$ and parameters controlling the signal strength, e.g. $\varepsilon_{m}$ and $P_{t,{\mathcal{G}}_{t}}$. With this observation, we can reject $H_{0}$ if $s$ is smaller than some critical value $c^{\prime}$ satisfying $\mathbb{E}_{1}s\leq c^{\prime}\leq\mathbb{E}_{0}s$. We call this test the sum test, and it can reliably detect the alternative hypothesis under slightly stronger assumptions than those stated in Theorem 3.4, as stated in the next proposition.
@@
    Let $q+p<1/2$ and $P_{t,{\mathcal{G}}_{t}}\leq 1-m^{-q}$ for all $t=1,...,m$. For the sum test with the critical value $c^{\prime}:=\mathbb{E}_{0}s-m^{1-(p+q)}/4$, the sum of type I and type II errors tends to 0 as $m\to\infty$.
@@
Higher Criticism.
In the literature on sparse mixture detection, higher criticism (HC) (Donoho and Jin, 2004, 2015; Tony Cai et al., 2011) is recognized as a powerful non-parametric method: in cases where the likelihood ratio test asymptotically separates the hypotheses, HC achieves the same separation, a property known as “adaptive optimality”. This insight is also relevant in our context. HC proceeds as follows: for $t=1,\dots,m$, we begin by ordering the $\zeta_{t}$ values in ascending order: $\zeta_{(1)}<\zeta_{(2)}<...<\zeta_{(m)}.$ Then we define the test statistic as:
@@
Since the statistic $\mathrm{HC}^{*}_{m}$ has a heavy-tailed distribution under the null hypotheses, and stringent tests based on heavy-tailed null distributions can lead to lower power in small samples, Donoho and Jin (2004, 2015) suggests a modified HC statistic involving the maximum over all $\zeta_{t}$’s greater than or equal to $1/m$, which is $\mathrm{HC}^{+}_{m}=\max_{1\leq t\leq m,\zeta_{(t)}\geq 1/m}\mathrm{HC}_{m,t}$ to enhance test power. We incorporate this modification in our experiments in subsequent sections. Additionally, we aim to select a critical value that controls type I errors at a specific significance level, such as $\alpha=0.01$. Following the recommendation of Tony Cai et al. (2011), a practical approach for choosing the critical value involves simulating the $\mathrm{HC}^{+}_{m}$ scores under the null hypothesis multiple times and using the top $\alpha$-percentile from the resulting empirical histogram. Given the slow weak convergence of the HC statistic, this method typically yields more accurate critical values than the theoretical consideration $\sqrt{(2+\delta)\log\log m}$.
Practical Considerations.
Lastly, we emphasize that assuming $P_{t,{\mathcal{G}}_{t}}$ is bounded away from 1 is not necessarily practical for language models: the probability of sampling a token from the green list can be arbitrarily close to 1, or even equal to 1, especially when some parts of the texts are deterministic. Nonetheless, previous results can be reinterpreted in a way that reflects the practical situation for language models: notice that if for a green token, the probability sum of its corresponding green list $P_{t,{\mathcal{G}}_{t}}$ is arbitrarily close to 1, then the corresponding $\zeta_{t}$ is sampled from a uniform distribution that is very close to $U[0,1]$. We can treat this $\zeta_{t}$ as if its corresponding token has been modified, i.e., this $\zeta_{t}$ does not belong to the set of $\varepsilon_{m}$ fraction of the $\zeta_{t}$’s that contribute to the watermark signal. In other words, Proposition 3.5 and Theorem 3.6 essentially says that as long as a sufficient fraction of the green tokens are sampled from the language model when $P_{t,{\mathcal{G}}_{t}}$ is bounded away from 1, the watermark signal can reliably be detected by the test based on the sum test or HC.
@@
$\zeta^{\prime}_{t}:=\zeta_{t}\mathds{1}_{\{w_{t}\text{ is green}\}}+(1-\zeta_{t})\mathds{1}_{\{w_{t}\text{ is red}\}},~{}s:=\sum_{t=1}^{n}\zeta^{\prime}_{t},~{}\text{and}~{}\mathrm{HC}^{*}_{n}:=\max_{1\leq t\leq n}\sqrt{n}\cdot\left(\frac{t/n-\zeta^{\prime}_{(t)}}{\zeta^{\prime}_{(t)}(1-\zeta^{\prime}_{(t)})}\right),$
@@ 3.3 Watermark Key
During detection, we select tokens carefully to compute the test statistics. While repeated $k$-grams generate the same seed, scoring only unique $k$-grams would be too restrictive, especially for stochastic decoders, where the same $k$-gram context may yield different next tokens. Instead, we score tokens based on the uniqueness of the full $(k+1)$-tuple, comprising the watermark context ($k$-gram) and current token. This approach prevents double-counting of identical subsequences while capturing valid watermark signals from stochastic decoding. As Fernandez et al. (2023) demonstrated, this scoring scheme improves true positive rates (TPR) compared to only scoring unique $k$-grams, while maintaining accurate false positive rate (FPR) estimates. This improvement results from capturing more valid watermark signals without redundant scoring.
While further refinement of watermark key usage and sharing is possible, these considerations are beyond the scope of this paper (Christ et al., 2023; Kirchenbauer et al., 2023b; Wu et al., 2023). Finally, we note that a single pre-generated green list ${\mathcal{G}}$ could be reused for all $t=1,\dots,n$, instead of generating a new green list at each $t$ (Zhao et al., 2023). In theory, the unbiasedness of our decoding algorithm (Lemma 2.2) is independent of the specific green list, a property also observed in practice; details are in Appendix A.5.4.
@@ 3.4 Interpret Speculative Decoding as Post-processing
Speculative decoding can be viewed as a model of the interaction between a human user and a language model. In this context, the smaller draft model represents the language model’s text generation capabilities, while the larger target model represents the human’s attempt to refine and improve the generated text. Just as a human user may find most of the text generated by a language model satisfactory but still make targeted modifications to align the text more closely with their intended meaning, style, and coherence, the larger model in speculative decoding validates and potentially modifies the predictions made by the smaller model.
While previous studies employed i.i.d. random substitution attacks to modify watermarked tokens (Fernandez et al., 2023; Kuditipudi et al., 2023), the post-processing in the speculative decoding setup offers a more realistic simulation of the interaction between a human user and a language model. By embedding a watermark only on the smaller draft model, we can explore its robustness under targeted modifications aimed at improving text quality. This practical setup also highlights an important consideration: while users may modify generated text, a high-quality watermarked model should minimize the need for such refinements, as frequent modifications risk compromising the watermark and lead to user fatigue. This analysis complements our theoretical analysis of the watermarking scheme and provides insights into the practical implications of the watermarking technique. The details on the implementation of our speculative decoding setup can be found in Appendix A.4.
@@ 4 Simulation Studies
Figure 2: The detection boundaries $12p+q=$ 1 and $p+q=1/2$ as shown in Theorem 3.6 and Proposition 3.5. The blue shaded area under the line $p+q=1/2$ represents the detectable regime for the sum tests. The orange shaded area under the line $12p+q=$ 1 shows the regime which is detectable for HC but not for the sum test. The red dots represent the choices of the parameters $p$ and $q$ we investigate in this simulation experiment, allowing us to investigate the power of both tests in four different scenarios: (1) fix $p=0.2$ and vary $q$ in $\{0.2,0.3,0.4,0.5,0.6\}$, (2) fix $q=0.4$ and vary $p$ in $\{0.1,0.2,0.3\}$, (3) on the boundary $p+q=1/2$, and (4) on the boundary $12p+q=$ 1.
We conduct simulation studies in two distinct scenarios to explore the robust detection boundaries discussed in Section 3.2. The first scenario focuses on situations where each green token carrying the watermark signal exhibits a sufficient signal strength, i.e., we only require $P_{t,{\mathcal{G}}_{t}}\geq m^{-r}$. This setting allows us to examine when detection is impossible, as revealed by Theorem 3.3. In contrast, the second scenario deals with a weak signal condition, i.e., $P_{t,{\mathcal{G}}_{t}}=1-m^{-q}$. Here, we aim to assess the effectiveness of both the sum test and HC in detecting signals, as detailed in Proposition 3.5 and Theorem 3.6.
[…]
The First Regime: Strong Signal.
We fix the value of $r$ at 0.2 and test different values of $p\in\{0.25,0.5,0.75\}$. To generate random variables $\zeta_{t}$ under the alternative hypothesis, we first sample $P_{t,{\mathcal{G}}_{t}}$ uniformly from $[m^{-r},1]$, and then sample $\zeta_{t}$ from $U[0,P_{t,{\mathcal{G}}_{t}}]$. For context, $0.1100^{-0.2}=0.398,1000^{-0.2}=0.251,10000^{-0.2}=0.158,100000^{-0.2}=$ 0.1. The results are shown in Figure 1. When $p$ is small, a large fraction of the $\zeta_{t}$’s carry the watermark signal, and the power of both tests converges to 1 as $m$ increases. However, as $p$ becomes larger, the power of both tests decreases. Indeed, with $p=0.75$, $r=0.2$ makes the problem undetectable as $12p-r=1.2>$ 1, and the power of both tests vanishes, consistent with Theorem 3.3. Interestingly, the power of the sum test is higher than that of HC, especially when $m$ is smaller. This observation will recur in the second regime.
@@
The Second Regime: Weak Signal.
We will verify the two theoretical boundaries, illustrated by the $p+q=1/2$ and $12p+q=$ 1 as shown in Figure 2. The shaded area below the $p+q=1/2$ boundary represents the cases where the sum test asymptotically separates the two hypotheses. The shaded area under the line $12p+q=$ 1 denotes the cases where HC asymptotically separates the two hypotheses. Each dot in the figure denotes a problem instance with the $(p,q)$ parameters accordingly selected. These selections enable a comprehensive evaluation of both tests across four distinct scenarios. For all simulations, we set $P_{t,{\mathcal{G}}_{t}}=1-m^{-q}$ to maintain a consistently weak signal strength.
@@
Scenario 1: Fix $p$ and vary $q$. We set the parameter $p$ at 0.2 and vary the parameter $q\in\{0.2,0.3,0.4,0.5,0.6\}$. Essentially, we examine how well each test performs as we maintain the proportion of signals but reduce their strength. The findings are illustrated in Figure 3. When both $p$ and $q$ are small, the power of both tests tends to converge to 1. As we move towards the boundary line $p+q=1/2$, the power of the sum test hovers around 0.3, while the power of HC increases to 1. Past this boundary, the sum test shows diminishing power, whereas HC’s power continues to rise, albeit at a slower pace. Upon reaching the upper boundary $12p+q=$ 1, both tests’ power trends towards zero. Interestingly, in scenarios with smaller $m$, the sum test often outperforms HC, consistent with our observation in the first regime. This suggests the sum test may be more suitable for analysis involving shorter texts.
@@
Scenario 2: Fix $q$ and vary $p$. We set the parameter $q$ at 0.4 and adjust the parameter $p\in\{0.1,0.2,0.3,0.4\}$. This approach allows us to investigate the effectiveness of both statistical tests as we maintain the strength of the signal but reduce the proportion of signals. The results are shown in Figure 4. The observation is similar to that in Scenario 1. At the boundary line $p+q=1/2$, the power of the sum test hovers around 0.3, while the power of HC increases to 1. As we move towards the boundary $12p+q=$ 1, both tests’ power trends towards zero, consistent with Theorem 3.3.
Scenario 3: On the boundary $p+q=1/2$. We keep $p+q=1/2$ and vary $p$ in $\{0.1,0.2,0.3\}$. That is, we stay on the detection boundary of the sum test while adjusting both the proportion and the strength of the signal. The results are shown in Figure 5. Across all situations, the power of the sum test stays around 0.3, while the power of HC approaches 1 more slowly as $p$ increases. Notice that with increasing $p$, the proportion of signals decreases dramatically: for example, $0.1261000^{-0.1}=0.501,1000^{-0.2}=0.251,1000^{-0.3}=$ 0.126. This trend indicates that the effectiveness of HC is more sensitive to changes in the proportion of watermarked tokens (as a function of $p$) than to the strength of the signal (as a function of $q$).
@@
Figure 6: The histograms of the test statistics under the null and alternative hypotheses for the sum test and HC when $m=100$. The test statistics are computed from the simulation under both the null and alternative hypotheses 2000 times. The critical value is determined by the quantile of the simulated histogram of the test statistic under the null hypothesis. The rejection rate, or the power, of each test, is computed by computing, under the alternative hypothesis, the proportion of the test statistics that are on the other side of the critical value. See also Figure 7 in Appendix.

### v2

@@ 3.2 Detection
Detection as Hypothesis Testing. Let $m$ be the number of green tokens in $\tilde{w}_{1:n}$. For simplicity, we assume that the first $m$ tokens in $\tilde{w}_{1:n}$, namely $\tilde{w}_{1},...,\tilde{w}_{m}$, represent the green tokens. If no corruption was made to the generated text, we can reformulate the detection objective into the following hypothesis testing problem:
@@ 2. $H_{1}$:
Sparse Mixture Detection. In practice, the user may modify the generated text. @hen a token is substituted, the corresponding uniform random variable $\zeta$ would be independent of the variable whether the new token is from the green list. In this case, the statistical signal of the watermark becomes sparse, leading us to consider an alternative hypothesis where a fraction $\varepsilon_{m}$ of the $\zeta_{t}$’s still present the watermark signal. Thus, we propose the following:
@@ 1. $H_{1}^{(\mathrm{mix})}$:
Sum Test. Since the probability sum of the green tokens, $P_{t,{\mathcal{G}}_{t}}$’s, are unknown to the detector, the likelihood ratio test cannot be used in practice. We begin by considering a simple test based on the sum $s=\sum_{i=1}^{m}\zeta_{m}$. Under the null hypothesis, $\mathbb{E}_{0}s=m/2$, while under the alternative hypothesis, $\mathbb{E}_{1}s$ is smaller than $m/2$ by a factor depending on $m$ and parameters controlling the signal strength, e.g. $\varepsilon_{m}$ and $P_{t,{\mathcal{G}}_{t}}$. With this observation, we can reject $H_{0}$ if $s$ is smaller than some critical value $c^{\prime}$ satisfying $\mathbb{E}_{1}s\leq c^{\prime}\leq\mathbb{E}_{0}s$. We call this test the sum test, and it can reliably detect the alternative hypothesis under slightly stronger assumptions than those stated in Theorem 3.5, as stated in the next proposition.
@@
    Let $q+p<1/2$ and $P_{t,{\mathcal{G}}_{t}}\leq 1-m^{-q}$ for all $t=1,...,m$. For the sum test with critical value $c^{\prime}:=\mathbb{E}_{0}s-m^{1-(p+q)}/4$, the sum of type I and type II errors tends to 0 as $m\to\infty$.
@@
Higher Criticism. In the literature on sparse mixture detection, higher criticism (HC) (Donoho and Jin, 2004, 2015) is recognized as a powerful non-parametric method: in cases where the likelihood ratio test asymptotically separates the hypotheses, HC achieves the same separation, a property known as “adaptive optimality.” This insight is also relevant in our context. HC proceeds as follows: for $t=1,\dots,m$, we begin by ordering the $\zeta_{t}$ values in ascending order: $\zeta_{(1)}<\zeta_{(2)}<...<\zeta_{(m)}.$ Then we define the test statistic as:
@@
Since the statistic $\mathrm{HC}^{*}_{m}$ has a heavy-tailed distribution under the null hypotheses, and stringent tests based on heavy-tailed null distributions can lead to lower power in small samples, Donoho and Jin (2004, 2015) suggests a modified HC statistic involving the maximum over all $\zeta_{t}$’s greater than or equal to $1/m$, which is $\mathrm{HC}^{+}_{m}=\max_{1\leq t\leq m,\zeta_{(t)}\geq 1/m}\mathrm{HC}_{m,t}$ to enhance test power. We incorporate this modification in our experiments in subsequent sections. Additionally, we aim to select a critical value that controls type I errors at a specific significance level, such as $\alpha=0.01$. Following Tony Cai et al. (2011), a practical approach for choosing the critical value involves simulating the $\mathrm{HC}^{+}_{m}$ scores under the null hypothesis multiple times and using the top $\alpha$-percentile from the resulting empirical histogram. Given the slow weak convergence of the HC statistic, this method typically yields more accurate critical values than the theoretical consideration $\sqrt{(2+\delta)\log\log m}$.
Practical Considerations. Lastly, we emphasize that assuming $P_{t,{\mathcal{G}}_{t}}$ is bounded away from 1 is not necessarily practical for language models: the probability of sampling a token from the green list can be arbitrarily close to 1, or even equal to 1, especially when some parts of the texts are deterministic. Nonetheless, previous results can be reinterpreted in a way that reflects the practical situation for language models: notice that if for a green token, the probability sum of its corresponding green list $P_{t,{\mathcal{G}}_{t}}$ is arbitrarily close to 1, then the corresponding $\zeta_{t}$ is sampled from a uniform distribution that is very close to $U[0,1]$. We can treat this $\zeta_{t}$ as if its corresponding token has been modified, i.e., this $\zeta_{t}$ does not belong to the set of $\varepsilon_{m}$ fraction of the $\zeta_{t}$’s that contribute to the watermark signal. In other words, Proposition 3.6 and Theorem 3.7 essentially says that as long as a sufficient fraction of the green tokens are sampled from the language model when $P_{t,{\mathcal{G}}_{t}}$ is bounded away from 1, the watermark signal can reliably be detected by the test based on the sum test or HC.
@@
$\zeta^{\prime}_{t}:=\zeta_{t}\mathbb{1}_{\{w_{t}\text{ is green}\}}+(1-\zeta_{t})\mathbb{1}_{\{w_{t}\text{ is red}\}},~{}s:=\sum_{t=1}^{n}\zeta^{\prime}_{t},~{}\text{and}~{}\mathrm{HC}^{*}_{n}:=\max_{1\leq t\leq n}\sqrt{n}\cdot\left(\frac{t/n-\zeta^{\prime}_{(t)}}{\zeta^{\prime}_{(t)}(1-\zeta^{\prime}_{(t)})}\right),$
@@ 3.3 Watermark Key
During detection, we select tokens carefully to compute the test statistics. While repeated $k$-grams generate the same seed, scoring only unique $k$-grams would be too restrictive, especially for stochastic decoders, where the same $k$-gram context may yield different next tokens. Instead, we score tokens based on the uniqueness of the full $(k+1)$-tuple, comprising the watermark context ($k$-gram) and current token. This approach prevents double-counting of identical subsequences while capturing valid watermark signals from stochastic decoding. As Fernandez et al. (2023) demonstrated, this scoring scheme improves true positive rates compared to only scoring unique $k$-grams, while maintaining accurate false positive rate estimates. This improvement results from capturing more valid watermark signals without redundant scoring. While further refinement of watermark key usage and sharing is possible, these considerations are beyond the scope of this paper (Christ et al., 2024; Kirchenbauer et al., 2023b). Finally, we note that a single pre-generated green list ${\mathcal{G}}$ could be reused for all $t=1,\dots,n$, instead of generating a new green list at each $t$ (Zhao et al., 2023). In theory, the unbiasedness of our decoding algorithm (Lemma 2.2) is independent of the specific green list, a property also observed in practice; details are in the online Appendix.
@@ 3.4 Interpret Speculative Decoding as Post-processing
Speculative decoding can be viewed as a model of the interaction between a human user and a language model: the smaller draft model represents the language model’s text generation capabilities, while the larger target model represents the human’s attempt to refine and improve the generated text. Just as a human user may find most of the text generated by a language model satisfactory but still make targeted modifications to align the text more closely with their intended meaning, style, and coherence, the larger model in speculative decoding validates and modifies the speculations made by the smaller model.
While previous studies employed i.i.d. random substitution attacks to modify watermarked tokens (Fernandez et al., 2023; Kuditipudi et al., 2023), the post-processing in the speculative decoding setup offers a more realistic simulation of the interaction between a human user and a language model. By embedding a watermark only on the smaller draft model, we can explore its robustness under targeted modifications aimed at improving text quality. This practical setup also highlights an important consideration: while users may modify generated text, a high-quality watermarked model should minimize the need for such refinements, as frequent modifications risk compromising the watermark and lead to user fatigue. This analysis complements our theoretical analysis of the watermarking scheme and provides insights into the practical implications of the watermarking technique. The details on the implementation of our speculative decoding setup can be found in the online Appendix.
@@ 4 Simulation Studies
[…]
The First Regime: Strong Signal. We fix the value of $r$ at 0.2 and test different values of $p\in\{0.25,0.5,0.75\}$. To generate random variables $\zeta_{t}$ under the alternative hypothesis, we first sample $P_{t,{\mathcal{G}}_{t}}$ uniformly from $[m^{-r},1]$, and then sample $\zeta_{t}$ from $U[0,P_{t,{\mathcal{G}}_{t}}]$. For context, $0.1100^{-0.2}=0.398,1000^{-0.2}=0.251,10000^{-0.2}=0.158,100000^{-0.2}=$ 0.1. The results are shown in Figure 1. When $p$ is small, a large fraction of the $\zeta_{t}$’s carry the watermark signal, and the power of both tests converges to 1 as $m$ increases. However, as $p$ becomes larger, the power of both tests decreases. Indeed, with $p=0.75$, $r=0.2$ makes the problem undetectable as $12p-r=1.2>$ 1, and the power of both tests vanishes, consistent with Theorem 3.4. Interestingly, the power of the sum test is higher than that of HC, especially when $m$ is smaller. This observation will recur in the second regime.
@@
The Second Regime: Weak Signal. We will verify the two theoretical boundaries, illustrated by the $p+q=1/2$ and $12p+q=$ 1 as shown in Figure 2. The shaded area below the $p+q=1/2$ boundary represents the cases where the sum test asymptotically separates the two hypotheses. The shaded area under the line $12p+q=$ 1 denotes the cases where HC asymptotically separates the two hypotheses. Each dot in the figure denotes a problem instance with the $(p,q)$ parameters accordingly selected. These selections enable a comprehensive evaluation of both tests across four distinct scenarios. For all simulations, we set $P_{t,{\mathcal{G}}_{t}}=1-m^{-q}$ to maintain a consistently weak signal strength.
Figure 2: The detection boundaries $12p+q=$ 1 and $p+q=1/2$ as shown in Theorem 3.7 and Proposition 3.6. The blue shaded area under the line $p+q=1/2$ represents the detectable regime for the sum tests. The orange shaded area under the line $12p+q=$ 1 shows the regime which is detectable for HC but not for the sum test. The red dots represent the choices of the parameters $p$ and $q$ we investigate in this simulation experiment.
@@
Scenario 1: Fix $p$ and vary $q$. We set the parameter $p$ at 0.2 and vary the parameter $q\in\{0.2,0.3,0.4,0.5,0.6\}$. Essentially, we examine how well each test performs as we maintain the proportion of signals but reduce their strength. The findings are illustrated in Figure 3. When both $p$ and $q$ are small, the power of both tests tends to converge to 1. As we move towards the boundary line $p+q=1/2$, the power of the sum test hovers around 0.3, while the power of HC increases to 1. Past this boundary, the sum test shows diminishing power, whereas HC’s power continues to rise, albeit at a slower pace. Upon reaching the upper boundary $12p+q=$ 1, both tests’ power trends towards zero. Interestingly, in scenarios with smaller $m$, the sum test often outperforms HC, consistent with our observation in the first regime, hinting the sum test may be more suitable for analysis involving shorter texts.
@@
Scenario 2: Fix $q$ and vary $p$. We set the parameter $q$ at 0.4 and adjust the parameter $p\in\{0.1,0.2,0.3\}$. This approach allows us to investigate the effectiveness of both statistical tests as we maintain the strength of the signal but reduce the proportion of signals. The results are shown in Figure 4. The observation is similar to that in Scenario 1. At the boundary line $p+q=1/2$, the power of the sum test hovers around 0.3, while the power of HC increases to 1. As we move towards the boundary $12p+q=$ 1, both tests’ power trends towards zero, consistent with Theorem 3.4.
@@
Scenario 3: On the boundary $p+q=1/2$. We keep $p+q=1/2$ and vary $p$ in $\{0.1,0.2,0.3\}$. That is, we stay on the detection boundary of the sum test while adjusting both the proportion and the strength of the signal. The results are shown in Figure 5. Across all situations, the power of the sum test stays around 0.3, while the power of HC approaches 1 more slowly as $p$ increases. With increasing $p$, the proportion of signals decreases dramatically: e.g. $0.1261000^{-0.1}=0.501,1000^{-0.2}=0.251,1000^{-0.3}=$ 0.126. This indicates that the effectiveness of HC is more sensitive to changes in the proportion of watermarked tokens (as a function of $p$) than to the strength of the signal (as a function of $q$).
@@
Figure 6: The histograms of the test statistics under the null and alternative hypotheses for the sum test and HC when $m=100$. The test statistics are computed from simulation under both the null and alternative hypotheses 2000 times. The critical value is set by the quantile of the simulated histogram of the test statistic under the null hypothesis.